Simplify and express each of the following in exponential form:  (2^3 × 3^4 × 4)/ (3 × 32)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Factors of 32 = 2 × 2 × 2 × 2 × 2
= 2^5
Factors of 4 = 2 × 2
= 2^2

Then,
= (2^3× 3^4× 2^2)/ (3 × 2^5)
= (2^3 + 2× 3^4) / (3 × 2^5) ... [∵(a^m x a^n ) = a^m+n
= (2^5× 3^4) / (3 × 2^5)
=2^5-5 x 3^4-1 ... [∵a^m÷a^n = a^m-n]
=2^0× 3^3
= 1 × 3^3
= 3^3

Thanks